# Computable Scott sentences and the weak Whitehead problem for finitely presented groups

### v2

@@ 1. Introduction
In [21] Scott proved that every countable structure is characterized up to isomorphism by a sentence in $\mathfrak{L}_{\omega_{1},\omega}$. Recently, there has been considerable interest [2, 6, 8, 9, 10, 11, 15] in understanding how simple such a sentence can been chosen, in the sense of following stratification of $\mathfrak{L}_{\omega_{1},\omega}$ from computable model theory:
@@
There exists a Hopfian finitely presented group with decidable word problem (so a computable such group) but undecidable weak Whitehead problem.
At first we thought that a positive answer to Conjecture 1.7 was known but correspondence with experts and a careful reading of the literature on the subject convinced us otherwise. There are two algorithmic problems close to this problem: the above-mentioned Whitehead problem and the generation problem. The latter problem asks about the decidability of determining if a given tuple of elements from the group generates the whole group. It is known that in $F_{2}\times F_{2}$ the Whitehead problem is decidable but the generation problem is not (cf. [16, Theorem 4.4]). The difficulty about finding an example as in Conjecture 1.7 is that both the decidability of the Whitehead problem and the decidability of the generation problem imply the decidability of the weak Whitehead problem, and so establishing undecidability of this latter problem is a harder task.
@@ 3. Groups defined by their types
It suffices to show that hyperbolic groups are finitely presented and Hopfian. The fact that they are finitely presented is well-known. Concerning Hopfianity, see e.g. [24]. ∎

### v3

@@ 1. Introduction
In [21] Scott proved that every countable structure is characterized up to isomorphism by a sentence in $\mathfrak{L}_{\omega_{1},\omega}$. Recently, there has been considerable interest [2, 6, 8, 9, 10, 11, 15] in understanding how simple such a sentence can been chosen, in the sense of the following stratification of $\mathfrak{L}_{\omega_{1},\omega}$ from computable model theory:
@@
There exists a Hopfian finitely presented group with decidable word problem (so a computable group) but undecidable weak Whitehead problem.
At first we thought that an affirmative answer to Conjecture 1.7 was known but correspondence with experts and a careful reading of the literature on the subject convinced us otherwise. There are two algorithmic problems close to this problem: the above-mentioned Whitehead problem and the generation problem. The latter problem asks about the decidability of determining if a given tuple of elements from the group generates the whole group. It is known that in $F_{2}\times F_{2}$ the Whitehead problem is decidable but the generation problem is not (cf. [16, Theorem 4.4]). The difficulty about finding an example as in Conjecture 1.7 is that both the decidability of the Whitehead problem and the decidability of the generation problem imply the decidability of the weak Whitehead problem, and so establishing undecidability of this latter problem is a harder task.
@@ 3. Groups defined by their types
It suffices to show that hyperbolic groups are finitely presented and Hopfian. The fact that they are finitely presented is well-known. Concerning being Hopfian, see e.g. [24]. ∎